OpenStudy (anonymous):

Consider the compound Sn(CO3)2 and answer the following questions. What is the name of this compound? What is the molar mass of this compound? Explain how you calculated this value THIS IS MY LAST QUESTION ON MY 45 QUESTION EXAM CAN SOME ONE PLEASE HELP..

OpenStudy (callisto):

Sn = Tin CO3 = Carbonate Since it's (CO3)2 , tin is of a charge of 2+ So, Sn(CO3)2 = Tin (II) carbonate From the periodic table Sn = 118.7 C = 12.0 O = 16.0 Molar mass of Sn(CO3)2 = 118.7 + ( 12.0 + 16.0 x 3 ) x 2 = ... ?